a) i) Suppose H0: "\mu" = "\mu" 0 is rejected in favour of H1 : "\mu" != "\mu" 0 at "\alpha" = 0.05 level of signifi cance. Would H0 necessarily be rejected at the "\alpha" = 0.01 level of signi cance? Explain.

ii) Suppose H0: "\mu" = "\mu" 0 is rejected in favour of H1 : "\mu" != "\mu" 0 at "\alpha" = 0.01 level of signfi cance. Would H0 necessarily be rejected at the "\alpha" = 0.05 level of signifi cance? Explain.


iii) If H0: "\mu" = "\mu" 0 is rejected in favour of H0: "\mu" > "\mu" 0, will it necessarily be rejected in favour of H1 : "\mu" != "\mu" 0 ? Assume that "\alpha" remains the same.




1
Expert's answer
2022-02-16T15:31:26-0500

"a)"

No,

When "H_0" is rejected at "\alpha=0.05", it implies that "p-value\lt \alpha=0.05". Changing the level of significance to "\alpha=0.01" would not necessarily lead to rejection of "H_0" since "0.01\lt p-value\lt 0.05".


"b)"

Yes,

When "H_0" is rejected at "\alpha=0.01", it implies that "p-value\lt \alpha=0.01". Changing the level of significance to "\alpha=0.05" would lead to rejection of "H_0" also since "p-value\lt 0.01\lt 0.05".


"c)"

No,

Let "p-value 1" be the p-value for an upper tailed test and "p-value2" be the p-value of a two tailed test.

When "H_0" is rejected at a given "\alpha" for an upper tailed test it shows that "p-value 1\lt \alpha".

The relationship between the p-value for these tests is,

"p-value1=1-{p-value 2\over 2}\implies p-value 2=2(1-(p-value 1))"

Clearly, the p-value of a two tailed test will be large since the upper tailed p-value was small. Thus, the null hypothesis would not be rejected.